Lynette is the CEO of publicly traded TTT Corporation and earns a salary of $200,000 in the current year. What is TTT Corporation’s after-tax cost of paying Lynette’s salary excluding FICA taxes?
Answer #1

Answer:-$130000

Assume ttt corporation marginal tax rate =35%

Before tax cost of salary =200000

After tax salary = salary*(1-tax rate)

=200000*(1-35%)=200000*65%=130000